### v2

@@ Appendix A Proofs of theorems
$Pr_{P^{n}}\left\{\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}\leq\frac{k\cdot n-1}{n}\right\}$
$=$ $Pr_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}-k\right)\leq-\frac{1}{\sqrt{n}}\right\}\rightarrow 0.5$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $-\frac{1}{\sqrt{n}}$ converges to zero. Denote $p_{R}=Pr\{R=1\}$. For the second term in (24), we have the following lower bound
$Pr_{P^{n}}\left\{\frac{1}{n}\sum_{i}R_{i}\geq\frac{k\cdot n}{n}\right\}$
$=$ $Pr_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}R_{i}-p_{R}\right)\geq\sqrt{n}\cdot\left(k-p_{R}\right)\right\}\rightarrow 1$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $\sqrt{n}\cdot\left(k-p_{R}\right)$ diverges to $\infty$- ∞ for $p_{R}>k$. We thus conclude
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}\geq 0.5$
@@
Similar argument shows $Pr_{P^{n}}\left\{\widehat{B}_{n}(\underline{t})>\widehat{B}_{n}(\widehat{t})\right\}$ has a limit of at least one half. We thus conclude
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)<B(\underline{t};P)\right\}\geq 0.5\Leftrightarrow\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}\geq 0.5$
@@
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}=\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}=0.5$
@@
$Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\leq 0\right\}$
$=$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\left\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}+\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}\leq 0\right\}$
$\leq$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}$
$\leq$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\frac{\min_{B(g;P)>k+c}\sqrt{n}\left(B(g;P)-k\right)}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $Pr_{P^{n}}\left\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{g\in\mathcal{G}}\Sigma^{B}(g,g)^{1/2}}\right\}$
@@
$Pr_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha}\ and\ B(g;P)>k\}=Pr_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $Pr_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $Pr_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
@@
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$=$ $\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\inf_{\text{ }g\in\mathcal{G}}\frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}=\alpha.$
@@
Proof of Corollary 3.2. The first part follows from Theorem 3.1, replacing $\alpha$ with $\alpha_{n}$ in its proof. Specifically, by construction we have
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{B(\hat{g}_{\alpha_{n}};P)>k\}\leq\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha_{n}}\ and\ B(g;P)>k\}\leq\alpha_{n}.$
@@
$\leq$ $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|+\widehat{W}_{n}(g_{P}^{\ast})-\widehat{W}_{n}(\hat{g}_{\alpha_{n}})$
@@
$Pr_{P^{n}}\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g_{P}^{\ast})-k\right)}{\widehat{\Sigma}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\leq c_{\alpha_{n}}\}\stackrel{{\scriptstyle a}}{{\sim}}\Phi\left(c_{\alpha_{n}}+\frac{\sqrt{n}\left(k-B(g_{P}^{\ast};P)\right)}{\Sigma_{P}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\right)\rightarrow 1$
for $B(g_{P}^{\ast};P)$ strictly below the threshold. Since $c_{\alpha_{n}}$ diverges to $\infty$- ∞ at a rate slower $\sqrt{n}$, the term in the parenthesis diverges to $\infty$ as $n\rightarrow\infty$. We then apply Lemma B.4 and B.6 in the Online Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1), which shows $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|$ converges to zero in probability. We then conclude $\hat{g}_{\alpha_{n}}$ is asymptotically welfare-efficient under distribution $P$. ∎
@@
$W(g_{P}^{\ast};P)=$ $\max_{B(g;P)\leq k}W(g;P)=\max_{B(g;P)\leq k}W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$
$\leq\max_{g\in\mathcal{G}}W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$
$=W(\widetilde{g}_{P};P)-\overline{\lambda}\cdot(B(\widetilde{g}_{P};P)-k)_{+}\leq W(\widetilde{g}_{P};P)$
@@
$B(\widetilde{g}_{P};P)-k\leq\frac{W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}.$
@@
Proof of Theorem 4.1. We first prove $\widehat{g}_{\text{tradeoff}}$ is asymptotically welfare-efficient. Denote the new objective function with $V(g;P)=W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$, whose maximizer is $\widetilde{g}_{P}$. Denote $\widehat{V}_{n}(g)=\widehat{W}_{n}(g)-\overline{\lambda}\cdot(\widehat{B}_{n}(g)-k)_{+}$ to be the sample-analog, whose maximizer is $\widehat{g}_{\text{tradeoff}}$. Apply uniform deviation bound to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $\widetilde{g}_{P}$, we have
@@
$=$ $2\sup_{g}\left|W(g;P)-\widehat{W}_{n}(g)-\overline{\lambda}\cdot(\max\{B(g;P)-k,0\}-\max\{\widehat{B}_{n}(g)-k,0\})\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\overline{\lambda}\cdot\sup_{g}\left|\max\{\widehat{B}_{n}(g)-k,0\}-\max\{B(g;P)-k,0\}\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\overline{\lambda}\cdot\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$
The last line uses the fact that $\left|\max\{a,0\}-\max\{b,0\}\right|\leq\left|a-b\right|$. Both terms, $\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|$ and $\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$ converge to zero in probability under Assumption 3.1. Specifically, Lemma B.4 and B.6 in the Online Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1) imply the uniform convergence in probability
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{W}_{n}(g)-W(g;P)\right|\rightarrow_{p}0,\ \sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{B}_{n}(g)-B(g;P)\right|_{p}\rightarrow 0.$
@@
Furthermore, we have $W(g_{P}^{\ast};P)\leq V(\widetilde{g}_{P};P)\leq W(\widetilde{g}_{P};P)$ as a by-product of Lemma 4.1 for any $P\in\mathcal{P}$. Putting these together, we have
@@
By the uniform deviation bound shown above,applied to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $g_{P}^{\ast}$, we have that uniformly with probability approaching one
@@
which is equivalent to $(B(\widehat{g}_{\text{tradeoff}};P)-k)_{+}\leq\frac{W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$. ∎
Online Appendix to Empirical Welfare Maximization with Constraints
This Online Appendix contains proofs of supporting lemmas and additional results stated in the paper.

### v3

@@ Appendix A Proofs of theorems
$\operatorname{Pr}_{P^{n}}\left\{\frac{1}{n}\sum_{i}C_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}\leq\frac{k\cdot n-1}{n}\right\}$
$=$ $\operatorname{Pr}_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}C_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}-k\right)\leq-\frac{1}{\sqrt{n}}\right\}\rightarrow 0.5$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $\frac{1}{\sqrt{n}}$ converges to zero. Denote $p_{C}=\operatorname{Pr}\{C=1\}$. For the second term in (14), we have the following lower bound
$\operatorname{Pr}_{P^{n}}\left\{\frac{1}{n}\sum_{i}C_{i}\geq\frac{k\cdot n}{n}\right\}$
$=$ $\operatorname{Pr}_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}C_{i}-p_{C}\right)\geq\sqrt{n}\cdot\left(k-p_{C}\right)\right\}\rightarrow 1$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $\sqrt{n}\cdot\left(k-p_{C}\right)$ diverges to -$\infty$ for $p_{C}>k$. We thus conclude
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}\geq 0.5$
@@
[…]
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{B(\widehat{t};P)<B(\underline{t};P)\right\}\geq 0.5\Leftrightarrow\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}\geq 0.5$
@@
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}=\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}=0.5$
@@
$\operatorname{Pr}_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\leq 0\right\}$
$=$ $\operatorname{Pr}_{P^{n}}\left\{\min_{B(g;P)>k+c}\left\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}+\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}\leq 0\right\}$
$\leq$ $\operatorname{Pr}_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}$
$\leq$ $\operatorname{Pr}_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\frac{\min_{B(g;P)>k+c}\sqrt{n}\left(B(g;P)-k\right)}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $\operatorname{Pr}_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $\operatorname{Pr}_{P^{n}}\left\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{g\in\mathcal{G}}\Sigma^{B}(g,g)^{1/2}}\right\}$
@@
$\operatorname{Pr}_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha}\text{ and }B(g;P)>k\}=\operatorname{Pr}_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $\operatorname{Pr}_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $\operatorname{Pr}_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
@@
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$=$ $\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{\inf_{\text{ }g\in\mathcal{G}}\frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}=\alpha.$
@@
Proof of Corollary 3.2 Uniformly asymptotic feasibility follows from Theorem 3.1, replacing $\alpha$ with $\alpha_{n}$ in its proof. Specifically, by construction we have
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{B(\hat{g}_{\alpha_{n}};P)>k\}\leq\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}\operatorname{Pr}_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha_{n}}\text{ and }B(g;P)>k\}\leq\alpha_{n}.$
@@
$\leq$ $\sup_{g\in\mathcal{G}}2|W(g;P)-\widehat{W}_{n}(g)|+\widehat{W}_{n}(g_{P}^{\ast})-\widehat{W}_{n}(\hat{g}_{\alpha_{n}})$
@@
$\operatorname{Pr}_{P^{n}}\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g_{P}^{\ast})-k\right)}{\widehat{\Sigma}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\leq c_{\alpha_{n}}\}\stackrel{{\scriptstyle a}}{{\sim}}\Phi\left(c_{\alpha_{n}}+\frac{\sqrt{n}\left(k-B(g_{P}^{\ast};P)\right)}{\Sigma_{P}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\right)\rightarrow 1$
for $B(g_{P}^{\ast};P)$ strictly below the threshold. Since $c_{\alpha_{n}}$ diverges to -$\infty$ at a rate slower $\sqrt{n}$, the term in the parenthesis diverges to $\infty$ as $n\rightarrow\infty$. We then apply Lemma B.4 and B.6 in the Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1), which shows $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|$ converges to zero in probability uniformly. We then conclude $\hat{g}_{\alpha_{n}}$ is asymptotically welfare-efficient uniformly under $\mathcal{P}$. ∎
Proof.
Proof of Corollary 3.3 Uniformly asymptotic feasibility follows exactly as in the Proof of Corollary 3.2. To prove uniformity in asymptotic welfare-efficiency, I slightly abuse notation by indexing the policy class by $\theta$ in the welfare and budget functions below. Since the policy class includes thresholding on continuously distributed $X$, and by assumption the conditional benefit and cost are positive, the welfare function $W(\theta;P)$ and the budget function $B(\theta;P)$ are both continuous and increases in $\theta$. That is, whenever $\theta\prec\theta^{\prime}$ where $\prec$ denotes element-wise inequality, we have $W(\theta;P)<W(\theta^{\prime};P)$ and $B(\theta;P)<B(\theta^{\prime};P).$ Therefore, at the constrained optimal threshold $\theta_{P}^{\ast}$, the constraint is binding $B(\theta_{P}^{\ast};P)=k.$ By the continuity and strict monotonicity of $W(\theta;P)$, for any $\epsilon>0$, there exists $\theta_{P,\epsilon}\prec\theta_{P}^{\ast}$ such that $W(\theta_{P,\epsilon};P)=W(\theta_{P}^{\ast};P)-\epsilon$.
Construct a sequence $b_{n}=-2c_{\alpha_{n}}/\sqrt{n}$, which decreases to zero since $c_{\alpha_{n}}$ diverges to -$\infty$ at a rate slower $\sqrt{n}$. Consider $B(\theta_{P}^{\ast};P)-B(\theta_{P,\epsilon};P)=k-B(\theta_{P,\epsilon};P)$, which is strictly positive for all $P\in\mathcal{P}$ because by assumption the conditional cost is bounded away from zero. Therefore there exists $N$ large enough such that for all $P\in\mathcal{P}$ we have
$k-B(\theta_{P,\epsilon};P)\geq b_{n}\cdot\Sigma_{P}^{B}(\theta_{P,\epsilon},\theta_{P,\epsilon})^{1/2}>0,\ \forall n\geq N.$ (15)
We decompose the welfare loss into
$W(\theta_{P}^{\ast};P)-W(\hat{\theta}_{\alpha_{n}};P)$
$=$ $W(\theta_{P}^{\ast};P)-W(\theta_{P,\epsilon};P)+W(\theta_{P,\epsilon};P)-W(\hat{\theta}_{\alpha_{n}};P)$
$=$ $\epsilon+W(\theta_{P,\epsilon};P)-\widehat{W}_{n}(\theta_{P,\epsilon})+\widehat{W}_{n}(\theta_{P,\epsilon})-\widehat{W}_{n}(\hat{\theta}_{\alpha_{n}})+\widehat{W}_{n}(\hat{\theta}_{\alpha_{n}})-W(\hat{\theta}_{\alpha_{n}};P)$
$\leq$ $\epsilon+\sup_{g\in\mathcal{G}}2|W(g;P)-\widehat{W}_{n}(g)|+\widehat{W}_{n}(\theta_{P,\epsilon})-\widehat{W}_{n}(\hat{\theta}_{\alpha_{n}})$
Under the event that $\theta_{P,\epsilon}\in\hat{\mathcal{G}}_{\alpha_{n}}$, we are guaranteed that the last term is non-positive. This happens with probability
$\operatorname{Pr}_{P^{n}}\{\frac{\sqrt{n}\left(\widehat{B}_{n}(\theta_{P,\epsilon})-k\right)}{\widehat{\Sigma}^{B}(\theta_{P,\epsilon},\theta_{P,\epsilon})^{1/2}}\leq c_{\alpha_{n}}\}\stackrel{{\scriptstyle a}}{{\sim}}\Phi\left(c_{\alpha_{n}}+\frac{\sqrt{n}\left(k-B(\theta_{P,\epsilon};P)\right)}{\Sigma_{P}^{B}(\theta_{P,\epsilon},\theta_{P,\epsilon})^{1/2}}\right).$
From (15), we have
$\lim_{n\rightarrow\infty}\inf_{P\in\mathcal{P}}\Phi\left(c_{\alpha_{n}}+\frac{\sqrt{n}\left(k-B(\theta_{P,\epsilon};P)\right)}{\Sigma_{P}^{B}(\theta_{P,\epsilon},\theta_{P,\epsilon})^{1/2}}\right)\geq\lim_{n\rightarrow\infty}\Phi\left(c_{\alpha_{n}}-2c_{\alpha_{n}}\right)=\lim_{n\rightarrow\infty}\Phi\left(-c_{\alpha_{n}}\right)=1$
as $c_{\alpha_{n}}$ diverges to -$\infty$ as $n\rightarrow\infty$. We then apply Lemma B.4 and B.6 in the Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1), which shows $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|$ converges to zero in probability uniformly over $P\in\mathcal{P}$. We then conclude $\hat{g}_{\alpha_{n}}$ is uniformly asymptotically welfare-efficient under $\mathcal{P}$. ∎
@@
$W(g_{P}^{\ast};P)=$ $\max_{B(g;P)\leq k}W(g;P)=\max_{B(g;P)\leq k}W(g;P)-\frac{\overline{\lambda}}{r}\cdot(B(g;P)-k)_{+}$
$\leq\max_{g\in\mathcal{G}}W(g;P)-\frac{\overline{\lambda}}{r}\cdot(B(g;P)-k)_{+}$
$=W(\widetilde{g}_{P};P)-\frac{\overline{\lambda}}{r}\cdot(B(\widetilde{g}_{P};P)-k)_{+}\leq W(\widetilde{g}_{P};P)$
@@
$B(\widetilde{g}_{P};P)-k\leq\frac{r\cdot(W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P))}{\overline{\lambda}}.$
@@
Proof of Lemma 4.2
Recall the new objective function with $V(g;P)=W(g;P)-\frac{\overline{\lambda}}{r}\cdot(B(g;P)-k)_{+}$, whose maximizer is $\widetilde{g}_{P}$. Recall $\widehat{V}_{n}(g)=\widehat{W}_{n}(g)-\frac{\overline{\lambda}}{r}\cdot(\widehat{B}_{n}(g)-k)_{+}$ is the sample-analog, whose maximizer is $\widehat{g}_{\text{tradeoff}}$. Apply uniform deviation bound to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $\widetilde{g}_{P}$, we have
@@
$=$ $2\sup_{g}\left|W(g;P)-\widehat{W}_{n}(g)-\frac{\overline{\lambda}}{r}\cdot(\max\{B(g;P)-k,0\}-\max\{\widehat{B}_{n}(g)-k,0\})\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\frac{\overline{\lambda}}{r}\cdot\sup_{g}\left|\max\{\widehat{B}_{n}(g)-k,0\}-\max\{B(g;P)-k,0\}\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\frac{\overline{\lambda}}{r}\cdot\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$
The last line uses the fact that $\left|\max\{a,0\}-\max\{b,0\}\right|\leq\left|a-b\right|$. Both terms, $\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|$ and $\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$ converge to zero in probability under Assumption 3.1. Specifically, Lemma B.4 and B.6 in the Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1) imply the uniform convergence in probability
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{W}_{n}(g)-W(g;P)\right|\rightarrow_{p}0,\ \sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{B}_{n}(g)-B(g;P)\right|\rightarrow_{p}0.$
@@
Proof.
Proof of Theorem 4.1.
By Lemma 4.1, we have $W(g_{P}^{\ast};P)\leq V(\widetilde{g}_{P};P)\leq W(\widetilde{g}_{P};P)$ for any $P\in\mathcal{P}$. Putting these together with Lemma 4.2, we have
@@
By the uniform deviation bound used in the proof for Lemma 4.2, applied to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $g_{P}^{\ast}$, we have that uniformly with probability approaching one
@@
which is equivalent to $(B(\widehat{g}_{\text{tradeoff}};P)-k)_{+}\leq\frac{r\cdot(W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P))}{\overline{\lambda}}$. ∎